Integrate (3-x^2)/(x^2+3)^2


1
Expert's answer
2021-11-16T11:36:38-0500

"\begin{aligned}
&\int \frac{3-x^{2}}{\left(x^{2}+3\right)^{2}} \mathrm{~d} x \\
&=-\int \frac{x^{2}-3}{\left(x^{2}+3\right)^{2}} \mathrm{~d} x \\
&=-\int\left(\frac{x^{2}+3}{\left(x^{2}+3\right)^{2}}-\frac{6}{\left(x^{2}+3\right)^{2}}\right) \mathrm{d} x \\
&=-\int\left(\frac{1}{x^{2}+3}-\frac{6}{\left(x^{2}+3\right)^{2}}\right) \mathrm{d} x \\
&=-\int \frac{1}{x^{2}+3} \mathrm{~d} x-6 \int \frac{1}{\left(x^{2}+3\right)^{2}} \mathrm{~d} x \\
&=-\frac{\arctan \left(\frac{x}{\sqrt{3}}\right)}{\sqrt{3}}+6\left[\frac{x}{6\left(x^{2}+3\right)}+\frac{1}{6} \int \frac{1}{x^{2}+3} \mathrm{~d} x\right] \\
&[\because \frac{1}{\left(\mathrm{a} x^{2}+\mathrm{b}\right)^{\mathrm{n}}} \mathrm{d} x=\frac{2 \mathrm{n}-3}{2 \mathrm{~b}(\mathrm{n}-1)} \int \frac{1}{\left(\mathrm{a} x^{2}+\mathrm{b}\right)^{\mathrm{n}-1}} \mathrm{~d} x+\frac{x}{2 \mathrm{~b}(\mathrm{n}-1)\left(\mathrm{a} x^{2}+\mathrm{b}\right)^{\mathrm{n}-1}} ]\\
&=-\frac{\arctan \left(\frac{x}{\sqrt{3}}\right)}{\sqrt{3}}+ \frac{x}{x^{2}+3}+\frac{\arctan \left(\frac{x}{\sqrt{3}}\right)}{\sqrt{3}}
=\frac{x}{x^{2}+3}
\end{aligned}"